INAF Pisces observing, 2011 November 28-29

Observer: C. Arcidiacono, G. Brusa, K. Boutsia, M. Fumana
Telescope Operator: D. Gonzalez Huerta
Support Astronomer: O. Kuhn
Instrument: Pisces

Summary

Thin clouds on the sky in the afternoon.
We open up early (4:00PM) to take narrow band flats with Pisces. Many flat frames are saturated due to cloud reflections. We obtained high and low count flats for Br-g, FeII, K, and some J.
At the beginning of the night the seeing was very good (~0.6'')

(JMH: This version is rolled back to match Rev 3 after it was accidentally overwritten with the following night's log.)

Data Summary

See the science log by Konstantina.Boutsia in ObsInaf20111129.

Details

23:04 Open up the dome - start NB flats
  • observe, flatBrg pisces.20111128.10723.fits - .10763.fits
  • counts varying between 5000counts and 15000counts - cloud passing??

23:35 observe,flatFeII,/no_tcs pisces.20111128.10764.fits - .10844.fits
  • some frames are saturated although we started with counts <19000 - still clouds?

00:15 observe,flatK,/no_tcs pisces.20111128.10845.fits - .10906..fits
  • pisces.20111128.10845.fits - 2sec
  • pisces.20111128.10846.fits - 7sec
  • the rest 8sec
  • there are saturated images 863-881

00:30 observe,flatJ,/no_tcs pisces.20111128.10907.fits - .10918.fits

00:42 starting NB flats for low counts
  • Brg pisces.20111128.10919.fits - .10938.fits
  • FeII pisces.20111128.10939.fits - .10958.fits
  • K pisces.20111128.10959.fits - .11013.fits
  • pisces.20111128.10993.fits - stray light

01:20 we finished with the flats - switched off ventilation/treehouse
  • pointing correction - collimation
  • at this point we also do 'pisces_set_archive' -> next image will be pisces.20111129.0001.fits

01:33 observe, IC10_preset (PA=-180, GS=8)
  • the active optics are collimated - but AO system started with 4micron rms
  • during the first preset CCD39 was blocked - we manually opened the loop and resend the preset
  • the second preset also failed - the loop seemed to be closed but the status of CCD39 was 'not live' - at some point it probably went in time out and the preset was cancelled - the slopes do not arrive from the BCU at the secondary
  • third preset failed again but in another point - the CCD39 is doing what it is supposed to do but nothing arrives at the secondary.
  • from AOS we switch OFF the WFS and start again : Wstop - Wstart
  • from AOS - turn ON
  • during this preset - the loop was closed - slopes started arriving at the secondary but there has been an 'acquire-ref FAIL' - preset cancelled
  • preset again - loop closed with 10 modes and then everything diverges and the mirror resets.

02:30 We restart everything! - AOS reboot

02:39 Call Alfio
  • what happened was: all the system had to be restarted and one subsystem (the arbitrator) had to be restared once more (software bug??)

02:48 problem seems to be solved by switcing OFF-ON the arbitrator - send preset again
  • irc_runao error - but presets seems to have failed before arriving at runao
  • send preset again
  • this time the preset was successful - the loop is closed and stable

03:08 after setting the 0,0 position we start dithering for IC10 in Ks 0008-0017
  • first dither image: pisces.20111129.0008.fits - exptime: 30sec
  • dither on target completed without error

03:17 preset cancelled when it was offsetting 145arcsec (sky image for this target)
  • the offset did not complete
JMH: The offset did complete at 03:16:41.
Nov 29 03:16:41 tcs2 LBT_IIF: IIF:Offset execution time: 12658 ms from IIF, handle 3011151
Nov 29 03:16:41 tcs2 LBT_IIF: irc:OffsetPointing command status: SUCCESS
JMH: But something happened to the secondary at 03:17:01.
Nov 29 03:17:01 tcs2 LBT_AOSR: [D] Updated variable adsecarb.R.LED:  1
Nov 29 03:17:01 tcs2 LBT_AOSR: [E] COILS DISABLED
Nov 29 03:17:01 tcs2 LBT_AOSR: [D] Updated variable adsecarb.R.FSM_STATE:  Failure
Nov 29 03:17:01 tcs2 LBT_AOSR: [E] Stopping WFS...
See Issue 3835.
  • offset and obtain sky-image manually - in order to do this we need to resend the preset - close loop - give offset for the sky position (instead of centering the target) and repeat dither
  • the preset was cancelled because the mirror was not yet set - repeat preset
  • the followed procedure (working script) is:
WORKING SCRIPT
# IC10_dither
# 1800sec Ks
header,PARTNER=INAF,OBSERVER='Bono',OBJECT=IC10
set_position
camera, EXP=30.0, FRAMES=2, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=1.0, ITER=1,IMAGE=SCIENCE
set_position
camera, EXP=30.0, FRAMES=2, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=5.0, ITER=1,IMAGE=SCIENCE
set_position
camera, EXP=30.0, FRAMES=2, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=5.0, ITER=1,IMAGE=SCIENCE
FAILING SCRIPT
# IC10_dither
# 1800sec Ks
header,PARTNER=INAF,OBSERVER='Bono',OBJECT=IC10
set_position
camera, EXP=30.0, FRAMES=2, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=1.0, ITER=1,IMAGE=SCIENCE
offset,X=145,Y=0,COO=DETXY145
camera, EXP=30.0, FRAMES=2, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=5.0, ITER=1,IMAGE=SCIENCE
offset,X=0,Y=0,COO=DETXY
camera, EXP=30.0, FRAMES=2, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=5.0, ITER=1,IMAGE=SCIENCE

03:55 preset cancelled - shell ripped for unknown reason 145
  • repeat preset - cancelled again because the shell was not flat yet...

03:57 the AO software is not responding
  • Adsec-Arbitrator is not responding - stop and restart
  • set the shell - success

04:03 send preset again - IC10_preset.obs
  • guider seeing 1.0'' - 1.2''

04:26 the preset was cancelled - shell ripped
  • it was skipping images and it failed - could it be that seeing was getting worse??
  • dimm seeing 1.6''
  • the AO ref star mag seemed to change - passing cirrus??

04:29 send preset again IC10
  • preset cancelled - it started skipping again

04:39 preset again - the flux is almost 1/3 than previously - probably the clouds are getting thicker

04:44 preset cancelled - too high forces sent at the secondary

04:48 moving to a bright target : T-Tau - send preset
  • 3 presets failed for various reasons (time out while trying to change bin, the system thought it was in closed loop but it was not...)

05:10 CCD39 went offline (not live)

05:23 send preset - T-Tau
  • dimm seeing 1.1'' - guider seeing 0.5''
  • CCD39 shows high noise
  • CCD39 Sensirion - 14.6C

05:31 after the preset we can see T-Tau - the loop is closed
  • we try to go on with the dither pattern

05:35 Maximum skip reached - AOS error
  • preset failed - try again

06:05 preset was successful - start dithering on T-Tau ==0077-

07:40 after a very long observe thru thicker and thinner clouds the loop failed

~07:41 restarted on same object

08:41 observed T Tauri with different PA to check if some features in the PISCES image are ghosts, they seem to be...

8:42 preset to NGC1560

9:40 the loop ended and the shell found at +40um to -32um PV di tip-tilt

9:48 AdSec Control GUI says Last Command: StopAo Not Processed

9:59 Loop is closed but AOS GUI shows WFS is OFF (see Snapshot)

10:24 loop failed again

10:31 active optics had huge offsets on both primary and secondary and the image in the acquisition camera looked completely out of focus (John, Andrew).

JMH: Somebody sent some huge and incorrect active optics corrections.
Nov 29 10:19:48 tcs3 LBT_PSFR: ACTOPT: Primary Zernikes 0 0 0 -19151.7 -2812.06 0 0 -17636.1 12193.7 -15175.2 0 0 -0 0 0 0 -0 0 -0 -0 0
Nov 29 10:19:48 tcs3 LBT_PSFR: ACTOPT: Secondary Zernikes 0 0 0 0 0 15420.1 17466 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Nov 29 10:20:41 tcs2 LBT_AOSR: [E] setZernikes(). PSF says: right PMC Actuator ChangeAllOffsets Command Failed because Bad new influence forces
The GCS WFS corrections reported in the log at 10:19:05 were OK, but then we saw these offloads from AOSR:
Nov 29 10:19:47 tcs3 LBT_PSFR: PSFR:setZernikes executed from AOSR, handle 180876
Nov 29 10:19:47 tcs2 LBT_AOSR: [I] HO Offload request. Z: 0 0 0 0 -19151.7 -2812.06 15420.1 17466 -17636.1 12193.7 -15175.2 0 0 -0 0 0 0 -0 0 -0 -0 0
See Issue 3836.

10:53 loop opened again with large tilt on the shell

10:56 doing one more preset AO

11:03 preset complete

11:15 shell ripped

11:24 preset to QSO0848: could not find the star.

11:35 the preset failed because after three active optics attempts it believed that the error was below 1000nm but it wasn't and the AO failed (Doug).

11:38 sent a second preset and now it has converged after > 3 attempts

11:40 on axis preset of the same object using an active preset, the issue seems to be that the combination of star brightness (faint) and residual aberration does not allow the acquire to work. Doing the active preset first we might be able to converge first with active and then one more time before the adaptive.

12:28 closed loop on the same object

-- KonstantinaBoutsia - 28 Nov 2011
Topic revision: r11 - 25 Jul 2014, DouglasFisher
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback